OpenStudy (anonymous):

The graph of a system of two linear equations has no solution. What is true about the lines? A. The lines are perpendicular. B. The lines have the same slope, but different intercepts. C. The lines have the same intercept, but different slopes. D. The lines are on top of each other.

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

@Johnbc

OpenStudy (anonymous):

@charlotte123

OpenStudy (anonymous):

actually john can you answer this one please

OpenStudy (anonymous):

At a school football game, an adult ticket costs $4 and a student ticket costs $1. A total of 250 people attend the game, earning the school $427. How many students attended the game?

OpenStudy (anonymous):

@Johnbc

OpenStudy (anonymous):

The lines must be have a similar X and Y value if they intersect the only way they do not is if they are parallel to each other. Parallel lines have the same slope.

OpenStudy (anonymous):

Create a system. \[250 people = Adults + Students\] \[(4 \times Adult) + (1 \times Students) = $427\]

OpenStudy (anonymous):

Solve the system for the value you are looking for.

OpenStudy (anonymous):

okay i did and it comes out to be 191

OpenStudy (anonymous):

Can not verify if there is no work shown.

OpenStudy (anonymous):

okkkkkkkkkkkkkkkkkkkkkkkkkkkkkk(: thanks anywasy